How many liters of a 14% alcohol solution must be mixed with 20L of a 50% solution to get a 30 % solution?

Answer by scott8148(6628)

x(14%)+20(50%)=(x+20)(30%) ... .14x+10=.3x+6 ... 4=.16x ... 25=x